# Beyond the Linear Separability Ceiling

### v1

@@ 7 Conclusion
This work demonstrates that the reasoning failures of VLMs in abstract visual tasks are not an issue of poor visual perception but a solvable misalignment in the subsequent reasoning process, revealing that these models possess far greater latent reasoning capabilities than their out-of-the-box performance suggests. Our analysis reveals that the linear separability ceiling, a measure of the discriminative power of a model’s embeddings to a linear classifier, is best understood not as an absolute performance limit, but as an observational ceiling that baseline models consistently fail to surpass. The linear reasoning bottleneck, defined by this ceiling, is therefore a prevalent but surmountable alignment issue. We show there is no single path to success; performance can be unlocked either by refining the model’s existing post-representation reasoning pathways to bypass the ceiling, or by restructuring its representations to make them more linearly separable and thus raise the ceiling itself.
Our key contributions include: (1) a novel diagnostic framework centered on the linear separability ceiling, which formalizes and quantifies the linear reasoning bottleneck in visual tasks; (2) an empirical demonstration that this bottleneck stems not from a fundamental capability deficit but from an alignment failure, proving that VLMs possess powerful dormant reasoning pathways, a finding supported by the success of postfix tuning which activates these pathways without altering the original visual representations; (3) a demonstration of the shifting performance bottleneck, showing that after successfully closing the linear reasoning bottleneck, the primary limitation shifts back to the quality of the visual representations themselves, making perception the next frontier for improvement; and (4) the discovery that objectives designed to improve representation quality can introduce structural brittleness, revealing an often overlooked trade-off between performance and structural robustness.
@@
This framework opens several avenues for future work, from developing training methods that avoid the trade-off between performance and structural robustness, to using the LSC as a live diagnostic during model development. Ultimately, our work provides both validated methods for enhancing VLM reasoning and novel insights into the representational properties of multimodal models, paving the way for more capable and interpretable AI.
Appendix A Baseline prompt definitions
@@
Strategies evaluated in Table 5 varied primarily in how images were presented relative to descriptive text, the position of the query image, and whether CoT reasoning was asked.
@@ 3. Provide your conclusion for the ‘query image‘ if it can be categorized as
Appendix B Single vs batched context
@@
We refer to these two distinct classification strategies as: Single context, a nearest-neighbor approach where the query image is classified based on its cosine similarity to the single closest example image vector from the combined positive and negative sets. This method is sensitive to the features of individual, and potentially outlier, examples; Batched context, a nearest-centroid approach selected for the LSC. The query image is classified based on its cosine similarity to a prototype vector (centroid) representing the entire set of positive or negative examples. This method is designed to capture the underlying abstract rule of the set. The results of this comparison are presented in Table 4.
Table 4: Comparison of classification accuracy for single vs. batched context.
@@
Appendix C Separability ceiling in VLMs
The Table 5 (next page) presents our baseline results, comparing each model’s generative accuracy to its LSC ($acc_{vision}$) and its final representation accuracy ($acc_{final}$).
While generative performance is not lower than their LSC in 84 out of 128 instances, only a handful of models manage to surpass this ceiling. This success is achieved through two distinct mechanisms. In 4 instances, models like Pixtral use non-linear representation refinement, where the model’s processing increases the linear separability of the embeddings ($acc_{final}>LSC$). More commonly, in 10 instances, models like Qwen-VL-72B employ a post-representation non-linear pathway, where generative performance exceeds the LSC despite a decrease in the linear separability of the final embeddings ($acc_{final}<LSC$).
A $\chi^{2}$ test highlights a complex relationship between embedding separability and generative output. The initial vision embeddings (LSC) show a statistically significant dependence on generative performance in 97 out of 128 instances, with only 2 of those being an inverse relationship. This suggests that the initial visual features are generally, and logically, aligned with the final output. However, the connection becomes more complicated after the model’s reasoning process. While the final embeddings have a significant dependence on the output in 103 instances, a surprising 29 of these are an inverse dependence. This underlines the disconnect between simple linear discriminability and the complex, non-linear reasoning happening within these models.
Table 5: Generative performance vs. the LSC. Superscripts on embedding scores denote significant (p<0.05) dependence (negative indicating inverse) with generative predictions (D=Direct, C=CoT).
Model Dataset Prompt Strategy Direct acc (%) CoT acc (%) Sim. acc (vision, %) Sim. acc (final, %)
@@ Appendix C Separability ceiling in VLMs
Interleaved query first $51.5\pm 3.5$ $55.5\pm 3.4$ $71.9\pm 3.1^{D,C}$ $56.5\pm 3.4^{D,C}$
@@
Labeled query first $55.1\pm 3.4$ $55.8\pm 3.4$ $71.9\pm 3.1^{D}$ $56.1\pm 3.4^{D,C}$
@@
Gemma3 27B OpenWorld Interleaved $91.8\pm 2.4$ $88.6\pm 2.8$ $88.6\pm 2.8^{D,C}$ $50.0^{-D,C}$
@@
HOI Interleaved $59.0\pm 3.4$ $62.1\pm 3.4$ $66.1\pm 3.3^{D,C}$ $57.6\pm 3.4^{-C}$
Interleaved query first $66.1\pm 3.3$ $62.0\pm 3.4$ $66.1\pm 3.3^{D,C}$ $58.6\pm 3.4^{-D,C}$
Labeled $54.0\pm 3.5$ $60.4\pm 3.7$ $66.1\pm 3.3^{D}$ $60.4\pm 3.4^{-D,-C}$
Labeled query first $66.5\pm 3.3$ $63.0\pm 3.5$ $66.1\pm 3.3^{D}$ $57.1\pm 3.4^{D,C}$
@@
Interleaved query first $77.1\pm 2.9$ $77.2\pm 2.9$ $72.1\pm 3.1^{D,C}$ $59.4\pm 3.4^{D,C}$
@@
Labeled query first $76.0\pm 3.0$ $77.5\pm 2.9$ $72.1\pm 3.1^{D,C}$ $56.2\pm 3.4^{D,C}$
Appendix D Hyperparameter details
LoRA and projector tuning employed a learning rate of $1e^{-4}$ and a batch size of 25. The LoRA experiments utilized a rank of $r=8$ and a scaling factor of $\alpha=8$. These hyperparameters are consistent with established practices for fine-tuning VLMs and training LoRA adapters (Hu et al., 2022; Liu et al., 2023; Microsoft, 2025).
@@ Appendix D Hyperparameter details
and then the the specific instruction word-by-word and nothing else!
@@
Hyperparameter search revealed that an aggressive configuration with a small batch size (2) and a high learning rate (0.01), which we refer to as the “fast” configuration, was highly effective for vision bias vector and postfix tuning. Prompt tuning uses the default learning rate of $1e^{-4}$ and a batch size of 25.
Appendix E Combined loss details
@@
Appendix F Combined loss hyperparameter search
[…]
Figure 4: Pixtral contrastive loss hyperparameter scan
Afterwards, target values were tested with different learning rates and since they proved stable, we picked these values. However, as seen with target value of 1.6 with learning rate 0.0001, the accuracy peaked quite early on. To address this early peaking and ensure a more controlled dynamic where the contrastive loss provides an initial strong signal that gradually yields to the NTP objective for final model refinement, we introduced a cosine learning schedule which has two separate schedules for $w_{n}$ and $w_{c}$, governed by the overall training progress. Let $p=\frac{\text{current\_step}+1}{\text{total\_steps}}$ represent the fraction of training completion, where current_step is the current training iteration (e.g., batch or epoch, 0-indexed) and total_steps is the total number of such iterations for the run.
The weight for the next-token prediction loss, $w_{n}$, is designed to smoothly increase, following the formula:
@@ Appendix F Combined loss hyperparameter search
Here, $C$ represents a constant scaling factor (0.4 for Phi and Gemma3 and 1.6 for Pixtral). This formulation for $w_{c}$ incorporates two dynamic components based on the training progress $p$. A cosine decay term, $\cos\left(\frac{\pi}{2}\cdot p\right)$, which smoothly transitions the base weight from $1$ down to 00. A linear ramp-up term, $\min\left(2\cdot p,1\right)$, which effectively scales the influence of the cosine term. This ramp-up is active during the first half of the training duration, remaining at $1$ for the second half. The product of these terms, further scaled by $C$, allows the contrastive loss to have an initial warm-up period and then a gradually diminishing influence as the training increasingly prioritizes the NTP objective towards its conclusion. Both prompt tuning and LoRA training was conducted using cosinel schedule. For prompt tuning progress was calculated using epochs and for LoRA, progress was calculated using batches.
Appendix G LoRA on vision encoder
@@
Table 6: Ablation of Vision Encoder LoRA. This table compares the final generative accuracy (%) when LoRA is applied to both the vision encoder and LLM versus only the LLM. The comparison is shown for both the standard next-token prediction ($\mathcal{L}_{\text{NT}}$) and the combined contrastive ($\mathcal{L}_{\text{combined}}$) objectives. The results show no statistically significant performance difference across these configurations.
@@
LoRA ($\mathcal{L}_{\text{NT}}$) $89.8\pm 2.7$ $89.8\pm 2.7$
LoRA ($\mathcal{L}_{\text{combined}}$) $96.2\pm 1.7$ $96.6\pm 1.6$
@@
LoRA ($\mathcal{L}_{\text{NT}}$) $77.4\pm 2.9$ $77.2\pm 2.9$
LoRA ($\mathcal{L}_{\text{combined}}$) $78.9\pm 2.8$ $78.2\pm 2.9$
@@
The results presented in Table 6 are remarkably consistent. In every tested scenario, there is no statistically significant performance difference between applying LoRA adapters to the LLM alone versus applying them to both the vision encoder and the LLM.
@@
Appendix H PEFT comparison
To understand how to best surpass the linear reasoning bottleneck, we compared two PEFT methods: prompt tuning, which seeks to activate latent abilities by finding an optimal input, and LoRA, which adapts the model’s core weights. We evaluated these methods using two distinct training objectives: the standard next-token prediction loss ($\mathcal{L}_{\text{NT}}$) and a combined objective that includes a contrastive loss ($\mathcal{L}_{\text{combined}}$) to explicitly improve representation quality. To evaluate structural robustness, we test performance on unseen prompt formats. The in-distribution (ID) evaluation uses the standard interleaved prompt structure seen during training. For the out-of-distribution (OOD) test, we use a labeled prompt structure that groups all images by category at the end of the prompt. This OOD setup challenges how well the optimization targets generalize beyond the syntactic structure on which they were trained.
Table 7: Comparison of PEFT methods. The in-distribution (ID) evaluation uses the standard interleaved prompt structure seen during training. For the out-of-distribution (OOD) test, we use a labeled prompt structure that groups all images by category at the end of the prompt. In the generative columns,bold indicates the performance is not significantly lower than the corresponding LSC, while an underline indicates the ceiling is surpassed. Superscripts on embedding classification scores indicate a statistically significant ($p<0.05$) dependence between the item-level predictions of the similarity-based classifier and a generative method: I for the ID method and O for the OOD method. Default configurations use a batch size of 25 and learning rate of $1e^{-4}$. The ’‘fast” configurations use a batch size of 2 and learning rate of 0.01.
@@ Appendix H PEFT comparison
Phi OpenWorld Direct baseline $59.0\pm 4.3$ $\mathbf{79.4\pm 3.5}$ $84.0\pm 3.2^{I}$ $76.4\pm 3.7^{-I}$ $78.2\pm 3.6$
Postfix tuning ($\mathcal{L}_{\text{NT}}$) $\underline{\mathbf{94.2\pm 2.0}}$ $\underline{\mathbf{90.2\pm 2.6}}$ $84.2\pm 3.2$ $77.4\pm 3.7^{I}$ $78.8\pm 3.6^{O}$
Prompt tuning ($\mathcal{L}_{\text{NT}}$) $\underline{\mathbf{90.4\pm 2.6}}$ $\mathbf{83.4\pm 3.3}$ $84.2\pm 3.2$ $74.6\pm 3.8^{I}$ $76.6\pm 3.7^{O}$
Prompt tuning ($\mathcal{L}_{\text{combined}}$) $\underline{\mathbf{94.4\pm 2.0}}$ $\mathbf{86.0\pm 3.0}$ $84.2\pm 3.2^{I}$ $94.2\pm 2.0^{I}$ $93.2\pm 2.2$
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{89.8\pm 2.7}$ $\mathbf{88.6\pm 2.8}$ $84.2\pm 3.2^{I}$ $74.0\pm 3.8$ $78.4\pm 3.6^{O}$
LoRA ($\mathcal{L}_{\text{combined}}$) $\underline{\mathbf{96.2\pm 1.7}}$ $\underline{\mathbf{94.2\pm 2.0}}$ $84.2\pm 3.2^{I}$ $93.2\pm 2.2^{I}$ $91.4\pm 2.5^{O}$

### v3

@@ 8 Conclusion
This work identifies a pervasive "alignment gap" in VLMs, where reasoning capabilities often lag behind the linear separability of their own visual perceptions on Bongard tasks. By introducing the Linear Separability Ceiling (LSC), we demonstrate that this bottleneck on baseline models is not stemming from poor perception but from weak top-down reasoning.
Our key contributions include:
    a novel diagnostic framework utilizing prototype vectors (centroids) to establish the LSC, which provides a geometric and statistical lens to formalize and quantify the alignment gap;
    the identification of two distinct pathways to surpass the LSC: enhancing linear separability or executing non-linear decision logic — where we show that the latter can be effectively activated via postfix tuning;
    a training methodology employing an auxiliary contrastive objective with a dual cosine schedule, which successfully balances the conflicting signals of representational structuring and next-token prediction;
    and empirical evidence linking the geometry of one-dimensionally linear representations to improved benchmark performance, supporting the hypothesis that globally consistent structures encourage transfer to out-of-distribution tasks.
@@
Our work provides both validated methods for enhancing VLM reasoning and novel insights into the representational properties of multimodal models, paving the way for more capable and interpretable AI.
9 LLM usage
The authors utilized a LLM to refine grammar and sentence structure, and to aid in discovering related work. The authors reviewed all generated content and assume full responsibility for this publication.
Appendix A LSC on VQA
To show that LSC is useful to disentangle perception from reasoning beyond Bongard tasks, we evaluate the Phi-3.5 model, its respective vision encoder (CLIP ViT-L/14) and its respective text encoder on multiple choice VQA tasks:
    Bongard-HOI and POPE (pope2025): Converted from binary to multiple-choice tasks operating on a single image. Each sample was restructured to contain one correct answer and three distractors (4 options total).
    A-OKVQA (aokvqa2022): A VQA task requiring world knowledge and commonsense reasoning.
    ScienceQA (scienceQA2022): A multimodal benchmark evaluating reasoning capabilities.
We established the LSC for these VQA tasks using the CLIP vision and text encoder. We compared this against a "random" baseline (simulating a model that always selects random answer), the standard Phi-3.5 baseline (zero-shot) and $\mathcal{L}_{NT}$ and $\mathcal{L}_{combined}$ fine-tuned models on Bongard HOI dataset. The results are summarized in Table A.1.
Table A.1: LSC analysis across VQA benchmarks. LSC (%) is the accuracy of a linear probe on the CLIP vision and text encoder. Random (%) denotes the accuracy when always selecting answer randomly. Gen. (%) is the performance of the Phi-3.5 model when prompted zero-shot without CoT. We compare the baseline Phi model to $\mathcal{L}_{NT}$ and $\mathcal{L}_{combined}$ fine-tuned models on Bongard HOI dataset.
Dataset LSC (%) Random (%) Baseline gen. (%) $\mathcal{L}_{NT}$ gen. (%) $\mathcal{L}_{combined}$ gen. (%)
HOI 78.75 25.04 78.89 67.57 54.53
POPE 62.53 25.00 79.65 71.34 68.40
A-OKVQA 61.14 24.54 74.34 64.72 55.73
ScienceQA 46.21 35.35 87.42 79.68 73.84
Conclusion. The scores between the LSC and the baseline model being identical for HOI dataset indicates the model is extracting all the linearly available information. For POPE, however, even when operating on the exact same COCO dataset images, the baseline model achieved results significantly higher than the LSC. This gap suggests two possibilities: it may be an artifact of image-concept memorization by Phi during its training, or it indicates that the necessary representations are inherently non-linear, requiring non-linear decision logic to be untangled. A-OKVQA and ScienceQA show the largest divergence, where the language-based abstract reasoning required exceeds the linear separability of the visual features alone. Furthermore, the effect of representational restructuring is most apparent on HOI and A-OKVQA, where the performance dropped significantly more with $\mathcal{L}_{combined}$ when compared to $\mathcal{L}_{NT}$. In contrast, POPE and ScienceQA are more resilient to this geometric shift. This suggests that the non-linear decision logic for these tasks relies less on representational geometry and more on memorized image-concept pairs, which remain robust to the imposed structural changes.
Appendix B Visualizing the manifold structure
Isomap is a non-linear dimensionality reduction technique that estimates the intrinsic geometry of a data manifold by preserving geodesic distances. While our PCA analysis demonstrated global linear separability, Figure A.1 reveals that the $\mathcal{L}_{combined}$ objective specifically promotes one-dimensionally linear representations, visualized here as distinct "rays" that persist across increasing neighborhood sizes ($k$). Furthermore, this geometric structure suggests that while semantic concepts are encoded as linear features (directions), the process of determining which linear axis is relevant for a specific query is context-dependent, necessitating the execution of non-linear decision logic to select the correct linear probe.
(a) Fine-tuned Phi with $\mathcal{L}_{NT}$ objective
(b) Fine-tuned Phi with $\mathcal{L}_{combined}$ objective
Figure A.1: Isomap projections of the positive concept centroids from the test set, colored by distance from the geometric center. (a) The $\mathcal{L}_{NT}$ trained model shows a manifold that loses structure as the neighborhood size ($k$) increases. (b) The model trained with $\mathcal{L}_{combined}$ reveals a manifold that is more robust as the neighborhood size ($k$) increases. The persistence of these "rays" at higher $k$ values indicates that semantic concepts are globally more consistent.
Appendix C Baseline prompt definitions
@@
Strategies evaluated in Table A.3 varied primarily in how images were presented relative to descriptive text, the position of the query image, and whether CoT reasoning was asked.
@@ 3. Provide your conclusion for the ‘query image‘ if it can be categorized as
Appendix D Metric extraction methodology
As illustrated in Figure A.2, performance metrics are derived from three distinct architectural stages. LSC is calculated by applying a nearest-centroid linear probe to the raw, non-contextualized embeddings directly from the vision encoder. The final representation accuracy (acc final) utilizes the same probe on the contextualized hidden states extracted from the last layer of the LLM. The generative accuracy (acc gen) is determined by evaluating the model’s standard next-token textual prediction against the ground truth.
Figure A.2: Flowchart illustrating the extraction points for visual embeddings and generative outputs within the VLM architecture.
Appendix E Single vs batched context
@@
We refer to these two distinct classification strategies as: Single context, a nearest-neighbor approach where the query image is classified based on its cosine similarity to the single closest example image vector from the combined positive and negative sets. This method is sensitive to the features of individual, and potentially outlier, examples; Batched context, a nearest-centroid approach selected for the LSC. The query image is classified based on its cosine similarity to a prototype vector (centroid) representing the entire set of positive or negative examples. This method is designed to capture the underlying abstract rule of the set. The results of this comparison are presented in Table A.2.
Table A.2: Comparison of classification accuracy for single vs. batched context.
@@
Appendix F Separability ceiling in VLMs
Table A.3: Generative performance vs. the LSC and the same linear probe accuracy on its final representations. Superscripts on embedding scores denote significant (p<0.05) dependence (negative indicating inverse) with generative predictions (D=Direct, C=CoT).
Model Dataset Prompt Strategy Direct acc (%) CoT acc (%) LSC linear probe (final, %)
@@ Appendix F Separability ceiling in VLMs
Interleaved query first $1.5\pm 3.5 55$$5.5\pm 3.4$ $71.9\pm 3.1^{D,C}$ $56.5\pm 3.4^{D,C}$
@@
Labeled query first $5.1\pm 3.4 55$$5.8\pm 3.4$ $71.9\pm 3.1^{D}$ $56.1\pm 3.4^{D,C}$
@@
Gemma3 27B OpenWorld Interleaved $91.8\pm 2.4$ $8.6\pm 2.8 88$$8.6\pm 2.8^{D,C}$ $50.0^{-D,C}$
@@
HOI Interleaved $59.0\pm 3.4$ $2.1\pm 3.4 66$$6.1\pm 3.3^{D,C}$ $57.6\pm 3.4^{-C}$
Interleaved query first $66.1\pm 3.3$ $2.0\pm 3.4 66$$6.1\pm 3.3^{D,C}$ $58.6\pm 3.4^{-D,C}$
Labeled $54.0\pm 3.5$ $0.4\pm 3.7 66$$6.1\pm 3.3^{D}$ $60.4\pm 3.4^{-D,-C}$
Labeled query first $66.5\pm 3.3$ $3.0\pm 3.5 66$$6.1\pm 3.3^{D}$ $57.1\pm 3.4^{D,C}$
@@
Interleaved query first $7.1\pm 2.9 77$$7.2\pm 2.9$ $72.1\pm 3.1^{D,C}$ $59.4\pm 3.4^{D,C}$
@@
Labeled query first $6.0\pm 3.0 77$$7.5\pm 2.9$ $72.1\pm 3.1^{D,C}$ $56.2\pm 3.4^{D,C}$
Appendix G Fine-tuning details
LoRA and projector tuning employed a learning rate of $1e^{-4}$ and a batch size of 25. Using LoRA we inject trainable matrices into the attention and MLP weights. LoRA experiments utilized a rank of $r=8$ and a scaling factor of $\alpha=8$. These hyperparameters are consistent with established practices for fine-tuning VLMs and training LoRA adapters (hu2022lora; liu2023visual; microsoft2025phi4mini).
@@ Appendix G Fine-tuning details
and then the specific instruction word-by-word and nothing else!
@@
Hyperparameter search revealed that an aggressive configuration with a small batch size (2) and a high learning rate (0.01), was highly effective for postfix tuning. Prompt tuning uses a learning rate of $1e^{-4}$ and a batch size of 25.
Appendix H Combined loss details
@@
The loss counterparts can be conflicting and pull weights in different directions, and to achieve consistent results, for Bongard task training we exercised a separate schedule for both parts.
Appendix I Combined loss hyperparameter search
To find hyperparameters for $\mathcal{L}_{\text{combined}}=w_{n}\mathcal{L}_{\text{NT}}+w_{c}\mathcal{L}_{\text{sim}}$, we fixed $w_{n}$ to 1 and scanned through $w_{c}$ values with different schedules. The constant schedule had the same $w_{c}$ throughout different epochs, which proved unstable for larger values. The linear schedule had a target value where every epoch the $w_{c}=\frac{1+\text{cur\_epoch}}{\text{total\_epochs}}\cdot\text{target\_value}$. Using linear schedule, scan was conducted over [0.1, 0.2, 0.4, 0.8, 1.6, 2.4, 3.2] target values. For Phi and Gemma3 the best value was 0.4 and for Pixtral best value was 1.6, which is shown on Figure A.3.
Figure A.3: Illustration of contrastive loss hyperparameter scan results for Pixtral. Plot displays model accuracy after every epoch (0-indexed) for selected hyperparameters.
Afterwards, target values were tested with different learning rates and since they proved stable, we picked these values. To ensure a more controlled dynamic where the contrastive loss provides an initial strong signal that gradually yields to the NTP objective for final model refinement, we introduced a cosine learning schedule which has two separate schedules for $w_{n}$ and $w_{c}$, governed by the overall training progress. This way we reduce conflicting learning signals and ensure similar results across different seeds. Let $p=\frac{\text{current\_step}+1}{\text{total\_steps}}$ represent the fraction of training completion, where current_step is the current training iteration (e.g., batch or epoch, 0-indexed) and total_steps is the total number of such iterations for the run. The weight for the next-token prediction loss, $w_{n}$, is designed to smoothly increase, following the formula:
@@ Appendix I Combined loss hyperparameter search
Here, $C$ represents a constant scaling factor (0.4 for Phi and Gemma3 and 1.6 for Pixtral). This formulation for $w_{c}$ incorporates two dynamic components based on the training progress $p$. A cosine decay term, $\cos\left(\frac{\pi}{2}\cdot p\right)$, which smoothly transitions the base weight from $1$ down to 0. A linear ramp-up term, $\min\left(2\cdot p,1\right)$, which effectively scales the influence of the cosine term. This ramp-up is active during the first half of the training duration, remaining at $1$ for the second half. The product of these terms, further scaled by $C$, allows the contrastive loss to have an initial warm-up period and then a gradually diminishing influence as the training increasingly prioritizes the NTP objective towards its conclusion. Both prompt tuning and LoRA training was conducted using dual cosine schedule. Progress for prompt tuning was calculated using epochs; for LoRA, progress was calculated using batches.
Appendix J LoRA on vision encoder
@@
Table A.4: Ablation of Vision Encoder LoRA. This table compares the final generative accuracy (%) when LoRA is applied to both the vision encoder and LLM versus only the LLM. The comparison is shown for both the standard next-token prediction ($\mathcal{L}_{\text{NT}}$) and the combined contrastive ($\mathcal{L}_{\text{combined}}$) objectives. The results show no statistically significant performance difference across these configurations.
@@
LoRA ($\mathcal{L}_{\text{NT}}$) $92.2\pm 2.4$ $92.4\pm 2.3$
LoRA ($\mathcal{L}_{\text{combined}}$) $95.6\pm 1.8$ $95.6\pm 1.8$
@@
LoRA ($\mathcal{L}_{\text{NT}}$) $8.6\pm 2.8 77$$7.5\pm 2.9$
LoRA ($\mathcal{L}_{\text{combined}}$) $9.2\pm 2.8 77$$7.9\pm 2.9$
@@
The results presented in Table A.4 are remarkably consistent. In every tested scenario, there is no statistically significant performance difference between applying LoRA adapters to the LLM alone versus applying them to both the vision encoder and the LLM.
@@
Appendix K PEFT comparison
To understand how best to tackle the alignment gap for this task, we compared two primary PEFT methods: prompt tuning, which seeks to activate latent abilities by finding an optimal input, and LoRA, which adapts the model’s core weights. We evaluated these methods using two distinct training objectives: the standard next-token prediction loss ($\mathcal{L}_{\text{NT}}$) and our combined objective ($\mathcal{L}_{\text{combined}}$), which adds a contrastive loss to explicitly improve representation quality. To further probe the limits of these methods, we tested both default and aggressive hyperparameter configurations for prompt-based tuning.
Our evaluation also measured structural robustness. The in-distribution (ID) test uses the standard interleaved prompt structure seen during training, while the out-of-distribution (OOD) test uses a labeled prompt structure that groups images by category, challenging the model’s ability to generalize beyond the training syntax.
Table A.5: Comparison of PEFT methods. In the generative columns, bold indicates the LSC is surpassed. Superscripts on embedding classification scores indicate a statistically significant (p<0.05p~<~0.05) dependence between the item-level predictions of the similarity-based classifier and a generative method: I for the ID method and O for the OOD method. By default we use batch size of 25 and learning rate of $1e^{-4}$. Postfix tuning uses an aggressive batch size of 2 and learning rate of 0.01.
@@
Phi OpenWorld Direct baseline $59.0\pm 4.3$ $79.4\pm 3.5$ $84.0\pm 3.2^{I}$ $76.4\pm 3.7^{-I}$ $78.2\pm 3.6$
Postfix tuning ($\mathcal{L}_{\text{NT}}$) $\mathbf{94.2\pm 2.0}$ $\mathbf{90.2\pm 2.6}$ $84.2\pm 3.2$ $77.4\pm 3.7^{I}$ $78.8\pm 3.6^{O}$
Prompt tuning ($\mathcal{L}_{\text{NT}}$) $\mathbf{90.4\pm 2.6}$ $83.4\pm 3.3$ $84.2\pm 3.2$ $74.6\pm 3.8^{I}$ $76.6\pm 3.7^{O}$
Prompt tuning ($\mathcal{L}_{\text{combined}}$) $\mathbf{94.4\pm 2.0}$ $86.0\pm 3.0$ $84.2\pm 3.2^{I}$ $94.2\pm 2.0^{I}$ $93.2\pm 2.2$
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{92.2\pm 2.4}$ $\mathbf{89.8\pm 2.7}$ $84.4\pm 3.2^{I}$ $74.8\pm 3.8$ $78.2\pm 3.6^{O}$
LoRA ($\mathcal{L}_{\text{combined}}$) $\mathbf{95.6\pm 1.8}$ $\mathbf{94.4\pm 2.0}$ $84.4\pm 3.2^{I}$ $93.8\pm 2.1^{I}$ $91.2\pm 2.5^{O}$